# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ 5 Theoretical Guarantees
In this section, we study the theoretical guarantees for the released network from our method. The very basic result is that our method gives differential privacy at the node level.
@@
Suppose Algorithm 1 is used to process the input network under the privacy budget $\varepsilon$. The release network $\tilde{A}^{11}$ is node $\varepsilon$-differentially private for each node $i\in[n]$.
Note that, even though we use the latent space model to motivate our method, the node DP guarantee does not rely on the latent space model. That means, even if the network data is not generated by latent space models, applying our method still protects privacy.
@@
5.1 Preservation of Network Properties: The Latent Distribution Consistency
As discussed, to study the preservation of network properties, we will assume that the true network data is generated from a latent space model.
@@ 5.1 Preservation of Network Properties: The Latent Distribution Consistency
We then introduce the following regularity assumptions for the model.
Assumption A2 (Latent distribution).
@@
  • If $d>1$, for each $l=2,\dots,d$:
    • The conditional CDF $F_{l|1:(l-1)}(x|u)$ is Lipschitz continuous in $(x,u)$ and strictly increasing in $x$.
    • The marginal density $f_{1:{(l-1)}}(u)$ is continuous, bounded and lower bounded by a positive constant.
@@
For each $1\leq j\leq d$, the estimation of the conditional CDF $F^{l|1:(l-1)}$ is constructed from $\{\hat{Z}_{i}\}_{i>n}$ by a kernel estimation using a bounded, Lipschitz continuous kernel $K:{\bf R}^{d-1}\to[0,\infty)$ (with a constant $L_{K}$) such that $\int_{{\bf R}^{d-1}}K(u)\,du=1$ and a properly chosen bandwidth $h=(\log m)^{-c}$ for a constant c>0c>0333This choice does not necessarily give the optimal rate in estimation. The optimal choice is hinged on the rate of A4 in a more complicated way, and the detailed study is not be the focus of our current work. We choose this rate for simple interpretations., defined as
$\widehat{F}^{\,l\mid 1:(l-1)}(x\,\big{|}\,x_{1},\dots,x_{l-1})=\frac{\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Z}_{n+i,l}\leq x\bigr{\}}\;K\!\Bigl{(}\tfrac{x_{1}-\hat{Z}_{i,1}}{h},\dots,\tfrac{x_{l-1}-\hat{Z}_{n+i,l-1}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{x_{1}-\hat{Z}_{n+i,1}}{h},\dots,\tfrac{x_{l-1}-\hat{Z}_{n+i,l-1}}{h}\Bigr{)}}.$ (9)
@@
Recall that we use $\{\tilde{Z}_{i}\}_{i\in[n]}$ to generate the released network $\tilde{A}^{11}$. Given the latent vectors, the generating mechanism is always independent Bernoulli following Definition 2. Therefore, to preserve network properties, the primary guarantee we need is that the resulting distribution of the $\tilde{Z}_{i}$’s should be roughly the same as the true distribution in a proper sense. Theorem 2 expresses this property.
@@
Suppose assumptions A1A4 hold. Let $\tilde{F}^{(i)}$ be the joint CDF of the random vector $\tilde{Z}_{i}$. For any $1\leq i\leq n$,
$\sup_{x\in{\bf R}^{d}}|\tilde{F}^{(i)}(x)-F(x)|\to 0$
in probability as $m\to\infty$.
On the one hand, we can say that the privatized $\tilde{Z}_{i}$ has the asymptotically correct distribution individually. On the other hand, due to the fact that $\{\tilde{Z}_{i}\}_{i\in[n]}$ cannot be treated as an i.i.d sample from $F$ collectively, the latent distribution consistency is for each individual node $i$, rather than a uniform convergence that holds simultaneously for all $i\in[n]$. Fortunately, although the uniform convergence is not achievable, the privatized latent vectors still give a similar collective behavior for the resulting network. Specifically, we now show that the empirical CDF of the privatized latent vectors is asymptotically consistent, that is, having behaviors similar to that of i.i.d true latent vectors.
@@
Suppose assumptions A1A4 hold. Let $\tilde{F}_{n}$ be the empirical CDF of the privatized latent vectors based on $\tilde{Z}_{i}\in{\bf R}^{d}$, $i=1,\cdots,n$ where
@@
$\sup_{x\in{\bf R}^{d}}|\tilde{F}_{n}(x)-F(x)|\xrightarrow{P}0.$
Assumption A4 for Theorems 2 and 3 requires that the latent vectors can be accurately recovered for all nodes $i\in[N]$. For nodes in the hold-out set, this is not difficult by using the standard network estimation methods. However, for nodes $i\in[n]$, such a requirement needs additional study of the node-wise estimation procedure. Next, we show that under both the inner product latent space model and the RDPG, the proposed estimation algorithm satisfies the requirement of A4.
@@
    The network is generated from the RDPG model and the latent vector set $\{\hat{Z}_{i}\}_{i\in[n]}$ is estimated by (7).
@@
$\max_{i\in[n]}\|{\hat{Z}_{i}-Z_{i}}\|\leq C\max(\delta_{m},\frac{\sqrt{\log{m}}+\sqrt{\log n}}{\sqrt{m}})$
@@
As specific examples, under both the inner product latent space model and RDPG, the assumption A4 in the latent distribution consistency results can be satisfied using commonly used estimation approaches.
@@
Suppose Assumptions A1A3 hold. In either of the following two cases:
@@
Subsequently, the established consistency of the latent distribution for the privatized latent vectors indicate correct recovery of many network properties.
5.2 Consistency of Privatized Network Moments
In this subsection, we show that, as a special application of the latent distribution consistency, the privatized network $\tilde{A}$ maintains the moments/motif counts of the original network.
Network moments are commonly used to measure various aspects of network structures for inference and comparison tasks (Cook, 1971; Milo et al., 2002; Bickel et al., 2011; Bhattacharyya and Bickel, 2015; Maugis et al., 2020; Levin and Levina, 2019; Zhang and Xia, 2022; Qi et al., 2024), as they allow comparison across networks of different sizes and node sets. More broadly, many other useful network statistics, though not directly expressed as network moments, can be written as functions of network moments of multiple motifs, such as the clustering coefficient. Therefore, the guarantee of network moment recovery can indicate the valid of many comparable analyses based on privatized data.
Recall that $\mathcal{V}(A)$ and $\mathcal{E}(A)$ are the node set and edge set of network $A$, respectively. Following Qi et al. (2024), a graph $R$ is a subgraph of $A$, written as $R\subset A$, if $\mathcal{V}(R)\subset\mathcal{V}(A)$ and $\mathcal{E}(R)\subset\mathcal{E}(A)$. Two graphs $R$ and $A$ are isomorphic, denoted by $R\cong A$, when there exists a bijective function $\phi$: $\mathcal{V}(R)\to\mathcal{V}(A)$ such that $(v_{i},v_{j})\in\mathcal{E}(R)$ if and only if edge $\phi(v_{i}),\phi(v_{j})]\in\mathcal{E}(A)$.
A motif refers to a (usually simple) graph, such as an edge (), a 2-star/V-shape (), a triangle (), or a 3-star (), which forms the building blocks of larger graphs. Here we denote a motif by $R$, with $|V(R)|=r$ representing the number of nodes. We focus exclusively on connected motifs. For a network $A$ and a motif $R$, the motif density of $R$ in $A$ is defined as the normalized number of subgraphs of $A$ that are isomorphic to $R$:
$X_{R}(A)=\big{|}\{S:S\subset A,S\cong R\}\big{|}/\binom{n}{r}=\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}{\mathbb{I}}(A_{[i_{1},\cdots,i_{r}]}\cong R)/\binom{n}{r}$, (10)
in which $A_{[i_{1},\cdots,i_{r}]}$ is the subgraph of $A$ induced by nodes $\{i_{1},\cdots,i_{r}\}$. We call $X_{R}(A)$ the network moment of $A$ with respect to motif $R$. For example, when $R$ is an edge, then $X_{R}(A)$ is the edge density of $A$. Network moments are summary statistics of the whole network structures. Intuitively, under the latent space model, they are determined by the collective behavior of the latent vectors. And therefore, preservation of latent vector distributions should also indicate the preservation of network moments. This is formally stated in the following result.
@@ 5.2 Consistency of Privatized Network Moments
$X_{R}(A^{11})-X_{R}(\tilde{A}^{11})\to 0$
in probability.
It is also trivial to see that the claim can be extended to any function of network motif counts such as the clustering coefficient. In the meantime, Theorem 5 provides crucial convenience in applications. For example, as demonstrated in Qi et al. (2024), one could compare multiple unmatchable networks by comparing the distribution of their moments. Theorem 5 indicates that we can conduct the same comparison inference using networks privatized by GRAND.
@@ 6 Simulation Experiments
In this section, we introduce simulation experiments to evaluate the performance of the proposed method. In the experiments, we fix the node DP budget, and assess the extent to which the privatized network can preserve the structural properties of the original network. We consider the two commonly used network models, inner-product latent space model and RDPG, as described in Section 3.3. For both methods, we generate adjacency matrices according to their respective models, and use them as the respective original datasets.
Experiment configuration
Under both models, we set certain key parameters as follows. We vary the released size and hold-out size to be $n=m\in\{2000,4000\}$, and network density $\rho\in\{0.025,0.05,0.1\}$. For example, for a network of size 2000, the expected average degree is 50 and 100, respectively. Also, we set the dimension $d\in\{3,6\}$ and the privacy budget $\varepsilon\in\{1,2,5,10\}$. Typically, $\varepsilon=1$ corresponds to a strong privacy protection in practice, while $\varepsilon=10$ gives much weaker protection. The latent distribution $F$ is a truncated Gaussian mixture model, as a generalization of the mechanism of Li et al. (2023a) for more heterogeneity.
Methods in comparison
In addition to our proposed method, we include two additional benchmark methods for comparison.
  1. 1.
    The first one is naively applying the Laplace mechanism of Dwork (2006) to $\{\hat{Z}_{i}\}_{i\in[n]}$ and then generating the network from the resulting latent vectors (“Laplace”). Following a similar argument as in Theorem 1, it also satisfies node DP, but offers no guarantee for structural preservation. The comparison with this benchmark illustrates how important it is to consider the latent distribution consistency in the design of the proposed method in practice.
  2. 2.
    We also consider a non-private method: We directly estimate the network models from the size $m$ network using a standard network estimation, rather than the node-wise estimation method. Specifically, we use the gradient descent model fitting of Ma et al. (2020); Li et al. (2023a) for the inner product model and the adjacency spectral embedding (Sussman et al., 2014) for the RDPG. After that, we generate a new size-$n$ network from the estimated model without introducing any privatization step as the result (in this case, we have $m=n$). We call this the “Hat” network. The “Hat” method corresponds to the scenario where we do not impose DP at all and only use a standard approach to estimate the network model and generate new data. Comparing the “Hat” network with the original network gives us a non-private benchmark. As discussed in Section 4, our model estimation before the privatization is essentially based on $m+1\approx m$ nodes. Therefore, the “Hat” estimate can be seen as a comparable one with our method when $\varepsilon\to\infty$. The comparison between our method and “Hat” can show the price of structural preservation we pay for the privacy.
Performance evaluation metrics
We measure the structural preservation by examining the distributions of five node-level statistics across the $n$ nodes in each released network, and compare these distributions with those of the original network. Specifically, for each local statistic, denote its value at node $i$ by $T_{i}(A^{11})$ on network $A^{11}$, and its value at node $i$ on the released network $\tilde{A}^{11}$ as $T_{i}(\tilde{A}^{11})$. We want to compare the distribution of $\{T_{i}(\tilde{A}^{11})\}_{1\leq i\leq n}$ with that of $\{T_{i}(A^{11})\}_{1\leq i\leq n}$. We focus on the following five local statistics as representative of network properties.
  1. 1.
    Node degree: $T_{i}(A)=\sum_{j}A_{ij}$.
  2. 2.
    V-shape count: $T_{i}(A)=\binom{\sum_{j}A_{ij}}{2}$. It measures how many V-shape motifs involve node $i$ as the center. To be consistent with the motif count definition, this also includes triangles.
  3. 3.
    Triangle count: $T_{i}(A)=\frac{1}{2}\sum_{j,k}A_{ij}A_{jk}A_{ki}$. It measures how many triangles involve node $i$.
  4. 4.
    Eigen-centrality: $T_{i}(A)=v(A)_{i}$ where $v(A)$ is the eigenvector corresponding to the largest eigenvalue of $A$. It is a popular measure of how central node $i$ is in the network based on the spectral structure.
  5. 5.
    Harmonic centrality: $T_{i}(A)=\frac{1}{\sum_{j\neq i}d_{A}(i,j)}$ where $d_{A}(i,j)$ is the geodesic distance between node $i$ and node $j$ in the network.
Therefore, for each pair of $A^{11}$ and $\tilde{A}^{11}$, we can visualize the resulting distributions of the five statistics and compare them. However, to aggregate the results in a meaningful way, we also have to introduce the metrics to measure the difference between a pair of distributions. In our study, we use the Wasserstein distance metric. For the highly skewed distributions (such as V-shape, triangle counts), directly calculating the distances becomes numerically unstable, so we applied a logarithm transformation to the data and evaluate the recovery of the log-transformed distributions. Under each configuration, we independently repeat the experiments 100 times and take the average distances as the final results.
6.1 Results under the inner product latent space model
@@
For the inner-product latent space model, we generalize the simulation setup of Li and Le (2023) to incorporate more heterogeneity. Following the model definition in Example 1, we sample latent vectors $X_{i}$’s from a mixture of truncated Gaussian distributions and generate $\alpha_{i}$’s from a uniform distribution. We then rescale all $\alpha_{i}$’s together so the resulting network from the model would have the desired density according to our previous configuration. We randomly sample half of the nodes to be the hold-out data, and privatize the network between the other half of the nodes.
Tables 1 and 2 give the distribution preservation errors for $n=2000$ and $n=4000$, respectively. In all experimental configurations, the distance is measured by the Wasserstein distance, and the network is generated from the inner product model. The results are summarized as follows.

### v3

@@ 5 Theoretical Guarantees
We now study the theoretical guarantees for the released network from our method. The very basic result is that our method gives differential privacy at the node level.
@@
Suppose Algorithm 1 is used to process the input network under the privacy budget $\varepsilon$. The resulting $\tilde{A}^{11}$ is node $\varepsilon$-differentially private.
Note that, even though we use the latent space model to motivate our method, the node DP guarantee does not rely on the latent space model. That means, even if the network data is not generated by latent space models, our method still protects privacy.
@@
5.1 Preservation of Network Properties
As discussed, to study the preservation of network properties, we will assume that the true network data is generated from a latent space model with regularity conditions.
@@ 5.1 Preservation of Network Properties
Assumption A2.
@@
  • If $d>1$, for each $l=2,\dots,d$: The conditional CDF $F_{l\,|\,1:(l-1)}(x\,|\,u)$ is Lipschitz continuous in $(x,u)$ and strictly increasing in $x$. The marginal density $f_{1:{(l-1)}}(u)$ is continuous, bounded and lower bounded by a positive constant.
@@
For each $1\leq l\leq d$, the estimation of the conditional CDF $F^{l\,|\,1:(l-1)}$ is constructed from $\{\hat{Z}_{i}\}_{i>n}$ by a kernel estimation using a bounded, Lipschitz continuous kernel $K:\mathbb{R}^{l-1}\to[0,\infty)$ (with a constant $L_{K}$) such that $\int_{\mathbb{R}^{l-1}}K(u)\,\mathrm{d}u=1$ and bandwidth $h=(\log m)^{-c}$ for a constant $c>0$, defined as
$\hat{F}^{l\,|\,1:(l-1)}(x\,|\,x_{1},\dots,x_{l-1})=\frac{\sum_{i=1}^{m}\mathbf{1}_{\{\hat{Z}_{n+i,l}\leq x\}}K\left(\tfrac{x_{1}-\hat{Z}_{n+i,1}}{h},\dots,\tfrac{x_{l-1}-\hat{Z}_{n+i,l-1}}{h}\right)}{\sum_{i=1}^{m}K\left(\tfrac{x_{1}-\hat{Z}_{n+i,1}}{h},\dots,\tfrac{x_{l-1}-\hat{Z}_{n+i,l-1}}{h}\right)}.$ (8)
@@
Recall that we use $\{\tilde{Z}_{i}\}_{i\in[n]}$ to generate the released network $\tilde{A}^{11}$. Given the latent vectors, the generating mechanism is always independent Bernoulli following Definition 2. Therefore, to preserve network properties, the primary guarantee we need is that the resulting distribution of the $\tilde{Z}_{i}$’s should be roughly the same as the true distribution in a proper sense.
@@
Suppose assumptions A1A4 hold. Let $\tilde{F}^{(i)}$ be the joint CDF of the random vector $\tilde{Z}_{i}$. For any $1\leq i\leq n$, as $m\to\infty$,
$\sup_{x\in\mathbb{R}^{d}}\big|\tilde{F}^{(i)}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0.$
On the one hand, we can say that the privatized $\tilde{Z}_{i}$ has the asymptotically correct distribution individually. On the other hand, due to the fact that $\{\tilde{Z}_{i}\}_{i\in[n]}$ cannot be treated as an i.i.d. sample from $F$ collectively, the latent distribution consistency is for each individual node $i$, rather than a uniform convergence that holds simultaneously for all $i\in[n]$. Fortunately, although the uniform convergence is not achievable, the privatized latent vectors still give a similar collective behavior for the resulting network. Specifically, we now show that the empirical CDF of the privatized latent vectors is asymptotically consistent.
@@
Suppose assumptions A1A4 hold. Let $\tilde{F}_{n}$ be the empirical CDF of the privatized latent vectors based on $\tilde{Z}_{i}\in\mathbb{R}^{d}$, $i\in[n]$ where
@@
$\sup_{x\in\mathbb{R}^{d}}\big|\tilde{F}_{n}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0.$
Assumption A4 for Theorems 2 and 3 requires that the latent vectors can be accurately recovered for all nodes $i\in[N]$. For nodes in the hold-out set, this is not difficult by using the standard network estimation methods. For nodes $i\in[n]$, such a requirement needs additional study of the node-wise estimation procedure. Next, we show that under both the inner product latent space model and the RDPG, the proposed estimation satisfies A4.
@@
    The network is generated from the RDPG model and the latent vector set $\{\hat{Z}_{i}\}_{i\in[n]}$ is estimated by the OLS linear regression.
@@
$\max_{i\in[n]}\|{\hat{Z}_{i}-Z_{i}}\|\leq C\max\left\{\delta_{m},\frac{\sqrt{\log{m}}+\sqrt{\log n}}{\sqrt{m}}\right\}$ (9)
@@
With the above results, under both models, we know the assumption A4 can be satisfied using commonly used estimation approaches.
@@
Suppose Assumptions A1A3 hold. In either of the following cases:
@@
Remark 6.
As mentioned before, we can adapt the main theoretical results to accommodate additional network sparsity parameters, but using different parameterizations for different models. For example, for the RDPG model, we can instead assume the connection probability to be $\rho_{n+m}Z_{i}^{\top}Z_{j}$ for some $\rho_{n+m}\to 0$ as $n,m\to\infty$, as used in Athreya et al. (2018). The other assumptions on about $Z_{i}$’s can remain the same. Theorems 2 and 3 still remain valid, as long as $m\rho_{m+n}\geq(\log{m})^{c^{\prime}}$ for a sufficiently large constant $c^{\prime}>0$ without significantly changing the standard density requirement (Athreya et al., 2018). The $\sqrt{m}$ term in (9) will be changed to $\sqrt{m\rho_{n+m}}$.
5.2 Preservation of Network Moments
In this subsection, we show that, as a result of the latent distribution consistency, the privatized $\tilde{A}$ maintains the moments/motif counts of the original network.
Network moments are commonly used to measure various aspects of network structures for inference tasks (Milo et al., 2002; Bickel et al., 2011; Bhattacharyya and Bickel, 2015; Maugis et al., 2020; Levin and Levina, 2019; Zhang and Xia, 2022; Qi et al., 2024), as they allow comparison across networks of different sizes and node sets. More broadly, many other useful network statistics, though not directly expressed as network moments, can be written as functions of network moments of multiple motifs, such as the clustering coefficient. Therefore, the guarantee of network moment recovery can indicate the valid of many downstream analyses under NDP.
Following Qi et al. (2024), a graph $R$ is a subgraph of $A$, written as $R\subset A$, if $\mathcal{V}(R)\subset\mathcal{V}(A)$ and $\mathcal{E}(R)\subset\mathcal{E}(A)$. Two graphs $R$ and $A$ are isomorphic, denoted by $R\cong A$, when there exists a bijective function $\phi$: $\mathcal{V}(R)\to\mathcal{V}(A)$ such that $(v_{i},v_{j})\in\mathcal{E}(R)$ if and only if edge $\phi(v_{i}),\phi(v_{j})]\in\mathcal{E}(A)$. A motif refers to a (usually simple) graph, such as an edge (), a 2-star/V-shape (), a triangle (), or a 3-star (), which forms the building blocks of larger graphs. Here we denote a motif by $R$, with $|V(R)|=r$ representing the number of nodes. We focus exclusively on connected motifs. For a network $A$ and a motif $R$, the motif density of $R$ in $A$ is defined as the normalized number of subgraphs of $A$ that are isomorphic to $R$:
$X_{R}(A)=\big|\{S:S\subset A,S\cong R\}\big|/\binom{n}{r}=\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}{\mathbb{I}}(A_{[i_{1},\ldots,i_{r}]}\cong R)/\binom{n}{r}$, (10)
in which $A_{[i_{1},\cdots,i_{r}]}$ is the subgraph of $A$ induced by nodes $\{i_{1},\cdots,i_{r}\}$. We call $X_{R}(A)$ the network moment of $A$ with respect to motif $R$. For example, when $R$ is an edge, then $X_{R}(A)$ is the edge density of $A$. Network moments are summary statistics of the whole network structures. Intuitively, under the latent space model, they are determined by the collective behavior of the latent vectors. And therefore, preservation of latent vector distributions should also indicate the preservation of network moments.
@@ 5.2 Preservation of Network Moments
$X_{R}(A^{11})-X_{R}(\tilde{A}^{11})\xrightarrow{\mathbb{P}}0.$
It is easy to see that the claim can be extended to any function of network motif counts such as the clustering coefficient. Theorem 5 provides crucial convenience in applications. For example, one can compare multiple unmatchable networks by comparing the distribution of their moments Qi et al. (2024), which can be done privately using GRAND-privatized networks.
@@ 6 Simulation Experiments
In this section, we introduce simulation experiments to evaluate the performance of the proposed method. In the experiments, we fix the node DP budget, and assess the extent to which the privatized network can preserve the structural properties of the original network. We consider the two commonly used network models, inner product latent space model and RDPG, as described in Section 3.2. For the inner product latent space model, our data generating method is a generalization of the mechanism of Li et al. (2023a) for more heterogeneity. We sample latent vectors $X_{i}$’s from a mixture of truncated Gaussian distributions and generate $\alpha_{i}$’s from a uniform distribution. We then rescale all $\alpha_{i}$’s together so the resulting network from the model would have the desired density according to our previous configuration. For the RDPG model, we sample latent vectors $Z_{1},\ldots,Z_{N}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\text{Uniform}(0,1)^{d}$. The $Z_{i}$’s are then rescaled to ensure the desired network density. We always randomly sample half of the nodes to be the hold-out data, and privatize the network between the other half of the nodes.
Experiment configuration. Under both models, we set certain key parameters as follows. We vary the released size and hold-out size to be $n=m\in\{2000,4000\}$, and network density $\rho\in\{0.025,0.05,0.1\}$. For example, for a network of size 2000, the expected average degree is 50, 100 and 200, respectively. Also, we set the dimension $d\in\{3,6\}$ and the privacy budget $\varepsilon\in\{1,2,5,10\}$. Typically, $\varepsilon=1$ corresponds to a strong privacy protection in practice, while $\varepsilon=10$ gives much weaker protection.
Methods in comparison. In addition to our proposed method, we include two additional benchmark methods for comparison. (a) The first one is naively applying the Laplace mechanism of Dwork (2006) to $\{\hat{Z}_{i}\}_{i\in[n]}$ and then generating the network from the resulting latent vectors (“Laplace”). Following a similar argument as in Theorem 1, it also satisfies node DP, but offers no guarantee for structural preservation. The comparison with this benchmark illustrates how important it is to consider the latent distribution consistency in the design of the proposed method in practice. (b) We also consider a non-private method: We directly estimate the network models from the size $m$ network using a standard network estimation, rather than the node-wise estimation method. Specifically, we use the gradient descent model fitting of Ma et al. (2020); Li et al. (2023a) for the inner product model and the adjacency spectral embedding (Sussman et al., 2014) for the RDPG. After that, we generate a new size-$n$ network from the estimated model without introducing any privatization step as the result (in this case, we have $m=n$). We call this the “Hat” network. The “Hat” method corresponds to the scenario where we do not impose DP at all and only use a standard approach to estimate the network model and generate new data. Comparing the “Hat” network with the original network gives us a non-private benchmark. As discussed in Section 4, our model estimation before the privatization is essentially based on $m+1\approx m$ nodes. Therefore, the “Hat” estimate can be seen as a comparable one with our method when $\varepsilon\to\infty$. The comparison between our method and “Hat” can show the price of structural preservation we pay for the privacy.
Performance evaluation metrics. We measure the structural preservation by examining the distributions of five node-level statistics across the $n$ nodes in each released network, and compare these distributions with those of the original network. Specifically, for each local statistic, denote its value at node $i$ by $T_{i}(A^{11})$ on network $A^{11}$, and its value at node $i$ on the released network $\tilde{A}^{11}$ as $T_{i}(\tilde{A}^{11})$. We want to compare the distribution of $\{T_{i}(\tilde{A}^{11})\}_{1\leq i\leq n}$ with that of $\{T_{i}(A^{11})\}_{1\leq i\leq n}$. We focus on the following five local statistics as representative of network properties. 1) Node degree: $T_{i}(A)=\sum_{j}A_{ij}$. 2) V-shape count: $T_{i}(A)=\binom{\sum_{j}A_{ij}}{2}$. It measures how many V-shape motifs involve node $i$ as the center. To be consistent with the motif count definition, this also includes triangles. 3) Triangle count: $T_{i}(A)=\frac{1}{2}\sum_{j,k}A_{ij}A_{jk}A_{ki}$. It measures how many triangles involve node $i$. 4) Eigen centrality: $T_{i}(A)=v(A)_{i}$ where $v(A)$ is the eigenvector corresponding to the largest eigenvalue of $A$. It is a popular measure of how central node $i$ is in the network based on the spectral structure. 5) Harmonic centrality: $T_{i}(A)=1/\sum_{j\neq i}d_{A}(i,j)$ where $d_{A}(i,j)$ is the geodesic distance between node $i$ and node $j$ in the network. With these definitions, for each pair of $A^{11}$ and $\tilde{A}^{11}$, we can visualize the resulting distributions of the five statistics and compare them. However, to aggregate the results in a meaningful way, we also have to introduce the metrics to measure the difference between a pair of distributions. In our study, we use the Wasserstein distance metric. For the highly skewed distributions (such as V-shape, triangle counts), directly calculating the distances becomes numerically unstable, so we applied a logarithm transformation to the data and evaluate the recovery of the log-transformed distributions.
We independently repeat each experiment 100 times and take the average distances as the final results. Due to the space limit, we present the results only for $\rho=0.05,n=4000$ in this section. Additional results for other configurations of $\rho=0.025,0.1,n=2000$ are included in Appendix I. The observed patterns are similar to this section. Tables 2 and 3 give the distribution preservation errors under both models and all privacy budget levels when $d=3$ and $6$. The results are summarized as follows.